# RESEARCH ARTICLE

### v1

@@ 3 DISCUSSION
This study introduces a novel cross-species and cross-modality framework for epileptic seizure detection using EEG data. Through multi-space alignment, we have developed a framework that not only addresses the fundamental challenges of EEG data variability across species and modalities, but also enhances the generalizability and accuracy of seizure detection.
@@ 3.1 Broad implications
To our knowledge, this is the first study that demonstrates the effectiveness of integrating heterogeneous data from different species and modalities to improve EEG-based seizure detection performance. However, the implications of this research extend beyond improving epilepsy diagnostics. The cross-species and cross-modality framework could revolutionize how neurological disorders are studied and treated, offering a model for developing diagnostic tools that are more broadly applicable. For clinical practice, this means diagnosis for populations where traditional approaches are incapable may be performed.
Our approach may also be generalizable to different brain-computer interface paradigms[33, 35], and suggests the possibility to combine data from different species/modalities to increase the amount of training data for large EEG models.
@@ 4 METHODS
We introduce our multi-space alignment approach below.
@@ 4.2 Input space alignment
Let the original EEG data be represented as $X\in\mathbb{R}^{N\times C\times T_{s}}$, where $N$ is the training batch size, $C$ the number of channels, and $Ts$ the number of time samples. The EEG data are first reshaped to $X\in\mathbb{R}^{N\times T_{s}\times C}$, and then input to the Transformer Encoder, which includes multi-head self-attention and a feedforward neural network. After this, a linear layer reduces the dimensionality for species with more channels. For instance, the human sEEG data in the CHSZ dataset have 18 channels, whereas the canine iEEG data in the Kaggle dataset have 16 channels; when transferring from the CHSZ dataset to the Kaggle dataset, ResizeNet reduces the human sEEG data to 16 channels to match the dimensionality of the canine iEEG data.
@@ 4.3 Feature space alignment
where $\mathcal{H}$ is a reproducing kernel Hilbert space with a feature mapping $\phi$. Extensions of MMD, e.g., joint MMD[36] and local MMD[37], can also be utilized in domain adaptation regularization.
@@ 4.4 Output space alignment
$L_{\mathrm{KD}}=\frac{1}{N}\sum_{i=1}^{N}\tau^{2}KL\left(\boldsymbol{p}_{r}^{\tau},\boldsymbol{p}_{s}^{\tau}\right),$ (5)
where $\tau$ is a relaxation parameter (referred as the temperature in[31]) to soften the output of the teacher network, and $p^{\tau}=\frac{\exp\left(z_{i}/\tau\right)}{\sum_{j}\exp\left(z_{j}/\tau\right)}$ is the softmax operation, with $z_{i}$ being the logit for the $i$-th class.
@@ 4.5 Multi-space alignment
To summarize, EA normalizes all data in the input space to a similar scale, ResizeNet projects higher dimensional data to a lower dimensionality, domain adaptation further aligns the feature distributions, and knowledge distillation finally aligns the model predictions.
@@ 4.6 Implementation details
Appendix A APPENDIX A: COMPARED APPROACHES
The proposed multi-space alignment method is presented in Figure A1.
@@ Appendix B APPENDIX B: DATASET INFORMATION
Table B1: Summary of the four epilepsy datasets

### v2

@@ 3 DISCUSSION
This study introduces a novel cross-species and cross-modality framework for epileptic seizure detection using EEG data. Through MSA, we have developed a framework that not only addresses the fundamental challenges of EEG data variability across species and modalities, but also enhances the generalizability and accuracy of seizure detection.
@@ 3.1 Broad implications
To our knowledge, this is the first study that demonstrates the effectiveness of integrating heterogeneous data from different species and modalities to improve EEG-based seizure detection performance. However, the implications of this research extend beyond improving epilepsy diagnostics. The cross-species and cross-modality framework could revolutionize how neurological disorders are studied and treated, offering a model for developing diagnostic tools that are more broadly applicable. For clinical practice, this means diagnosis for populations where traditional approaches are incapable may be performed. This pilot study offers valuable insights into the challenges and opportunities associated with integrating multi-species and multi-modality data. It lays the foundation for future efforts to collect large-scale EEG datasets, enabling the training of large brain models.
Our approach may also be generalizable to different brain-computer interface paradigms[43, 45], and suggests the possibility to combine data from different species/modalities to increase the amount of training data for large EEG models. The performance of the proposed ResizeNet+MSA approach on the seizure prediction task is presented in Table E2, demonstrating its effectiveness in this application.
@@ 4 METHODS
We introduce our MSA approach below.
@@ 4.2 Input space alignment
Let the original EEG data be represented as $X\in\mathbb{R}^{N\times C\times T_{s}}$, where $N$ is the training batch size, $C$ the number of channels, and $T_{s}$ the number of time samples. The EEG data are first reshaped to $X\in\mathbb{R}^{N\times T_{s}\times C}$, and then input to the Transformer Encoder, which includes multi-head self-attention and a feedforward neural network. After this, a linear layer reduces the dimensionality for species with more channels. For instance, the human sEEG data in the CHSZ dataset have 18 channels, whereas the canine iEEG data in the Kaggle dataset have 16 channels; when transferring from the CHSZ dataset to the Kaggle dataset, ResizeNet reduces the human sEEG data to 16 channels to match the dimensionality of the canine iEEG data.
@@ 4.3 Feature space alignment
where $\mathcal{H}$ is a reproducing kernel Hilbert space with a feature mapping $\phi$, and $\hat{X}_{t}$ is the ResizeNet transformation of ${X}_{t}$ defined in Eq. (3). Extensions of MMD, e.g., joint MMD[46] and local MMD[47], can also be utilized in domain adaptation regularization.
@@ 4.4 Output space alignment
$L_{\mathrm{KD}}=\frac{1}{N}\sum_{i=1}^{N}\tau^{2}KL\left(\boldsymbol{p}_{r},\boldsymbol{p}_{s}\right),$ (5)
where $\tau$ is a relaxation parameter (referred to as the temperature in[39]) to soften the output of the teacher network, and $p=\frac{\exp\left(z_{k}/\tau\right)}{\sum_{j}\exp\left(z_{j}/\tau\right)}$ is the softmax operation, with $z_{i}$ being the logit for the $i$-th class and $n$ the total number of classes.
@@ 4.5 Multi-space alignment
To summarize, EA normalizes the input data to a unified scale, ResizeNet reduces higher-dimensional data to a lower-dimensional representation, domain adaptation aligns feature distributions, and knowledge distillation ensures consistency in model predictions. Together, these alignments form a unified framework tailored to address the complexities of cross-species and cross-modality EEG analysis. By integrating innovations across the input, feature, and output spaces, the proposed approach achieves effective data normalization, robust feature representation, and consistent predictive performance. The MSA framework is illustrated in Figure A1.
@@ 4.6 Implementation details
Appendix A APPENDIX A: ALGORITHMS UNDER COMPARISON
The proposed MSA framework is illustrated in Figure A1. ResizeNet (Figure A1b) incorporates a Transformer encoder, a linear layer, and two reshape modules. MSA (Figure A1a) integrates alignments across input, feature, and output spaces for cross-species adaptation. Specifically, ResizeNet addresses input space discrepancies, domain adaptation aligns cross-species data in the feature space (Figure A1d), and knowledge distillation ensures output space alignment between species (Figure A1c). Details of the training and test set splits for the three experiment scenarios (within-species, unsupervised cross-species transfer, and semi-supervised cross-species transfer) are illustrated in Figure A2.
@@ Appendix A APPENDIX A: ALGORITHMS UNDER COMPARISON
Figure A2: Illustration of three experiment scenarios, taking the Human-to-Canine transfer task as an example.
@@
The classification accuracies (%) of the baseline methods and the proposed ResizeNet+ approaches are summarized in Table A1.
Table A1: Average classification accuracies (%) of unsupervised and semi-supervised cross-species transfer on Kaggle dataset. The best average accuracy of each task is marked in bold, and the second best by an underline.
Approach Canine-to-Human Human-to-Canine
0% 5% 10% 15% 20% Avg. 0% 5% 10% 15% 20% Avg.
Baseline Within / 81.77 83.13 86.31 86.72 84.48 / 85.33 87.61 90.39 91.73 88.77
Comb. 79.21 88.52 89.17 89.20 89.86 87.19 88.21 90.31 90.01 90.29 89.76 89.72
DAN 85.11*** 87.14 88.58 89.29 90.41 88.11 85.25 90.48 91.46* 92.90** 94.17** 90.85
JAN 85.03**** 86.32 87.33 88.40 89.66 87.35 85.11 89.41 91.17* 93.10*** 93.22* 90.40
SHOT 72.74 77.00 78.92 80.28 81.90 78.17 75.60 80.90 84.84 85.79 86.54 82.73
DSAN 82.89* 86.27 86.91 88.66 87.52 86.45 86.45 90.58 92.34** 93.30 94.16** 91.37
MCC 75.89 87.34 89.54 89.75 91.59 86.82 80.08 92.03 92.11** 93.84*** 93.85** 90.38
ResizeNet+ AT 82.92 89.56* 90.88** 92.14* 93.58** 89.82 89.28* 90.59 91.08 92.68* 93.66** 91.46
NST 85.02*** 87.49 88.23 89.01 89.70 87.89 88.35 91.39 91.66 91.38 91.40* 90.84
SP 84.38** 89.32 91.27* 92.24 94.06** 90.25 89.10 90.75 91.91** 92.59* 92.89* 91.45
RKD 80.22 89.14 91.07** 91.20 92.65** 88.86 88.04 90.96 91.94 91.37 92.93 91.05
PKT 83.04* 89.54 90.41 91.77 93.65** 89.68 89.18 90.75 92.21** 92.66* 93.77** 91.71
CC 85.19*** 89.84 91.53* 92.24* 93.46* 90.45 88.35 91.07 91.30 91.83 92.68* 91.05
MSA 85.55*** 90.29** 91.26** 92.34* 93.54** 90.60 89.58* 92.90* 93.72*** 94.80*** 95.05*** 93.21
  • ****: $p<0.0001$; ***: $p<0.001$; **: $p<0.01$; *: $p<0.05$.
@@ Appendix B APPENDIX B: DATASET INFORMATION
Table B1: Summary of the four epilepsy datasets.
@@ Appendix C APPENDIX C: DATA PREPROCESSING
Appendix D APPENDIX D: CROSS-DATASET HUMAN-TO-HUMAN TRANSFER EXPERIMENTS
Within-species cross-dataset experiments were conducted to validate the effectiveness of the proposed ResizeNet+MSA approach.
There were significant between-dataset discrepancies in this setting. For example, the CHSZ dataset includes 18 sEEG channels sampled at 400 Hz, whereas the Kaggle-Human dataset consists of 16-72 iEEG channels sampled at 500 Hz. Although these datasets originate from the same species, they differ significantly in acquisition methods, channel numbers and locations, and sampling rates. Based on these variations, we explored three experiment settings, as illustrated in Figure D1:
  1. 1.
    Cross-dataset, within-species (human-to-human) and within-modality transfer: In this scenario, models were trained on human sEEG data from the NICU or CHSZ dataset and tested on another human sEEG dataset. Since there were no channel differences between the NICU and CHSZ datasets, the proposed ResizeNet+MSA approach was not applied in this setting.
  2. 2.
    Cross-dataset, within-species (human-to-human) and cross-modality transfer: Models were trained on human iEEG or sEEG datasets and tested on human datasets of a different modality (e.g., iEEG to sEEG, or sEEG to iEEG). Input discrepancies exist in this setting, making alignment necessary.
  3. 3.
    Cross-dataset, multi-species [(canine+human)-to-human] and multi-modality transfer: Canine iEEG and human iEEG/sEEG datasets were combined as the training set, with CHSZ used as the test dataset. This setting involves both multi-species and multi-modality data, introducing significant input discrepancies.
Figure D1: Illustration of three cross-dataset transfer scenarios.
Table D1: Average unsupervised cross-dataset transfer AUCs (%) on NICU and CHSZ datasets. The best average performance of each task is marked in bold, and the second best by an underline.
[b] Approach Within-species Within-species Multi-species within-modality1 cross-modality2 multi-modality3 NICU CHSZ Kaggle (H) CHSZ NICU+Kaggle (C) NICU+Kaggle (C+H) -to- -to- -to- -to- -to- -to- CHSZ NICU CHSZ Kaggle (H) CHSZ CHSZ (sEEG (sEEG (iEEG (sEEG (iEEG+sEEG (iEEG+sEEG -to- -to- -to- -to- -to- -to- sEEG) sEEG) sEEG) iEEG) sEEG) sEEG) Baseline4 Source Only6 85.17±1.3 68.87±1.0 84.37±0.1 67.24±1.3 86.33±0.69 87.05±1.04 DAN 82.28±2.1 68.84±0.1 76.75±2.8 72.66±3.5 82.57±1.13 82.72±1.69 JAN 80.81±0.9 64.52±0.7 70.01±1.7 73.79±1.9 81.35±1.86 80.10±3.34 SHOT 80.03±1.1 68.82±0.2 72.75±3.5 53.92±1.5 74.44±4.08 80.86±1.00 DSAN 68.97±2.5 69.66±0.1 67.99±1.8 66.51±3.6 60.09±1.14 60.90±1.21 MCC 83.71±0.3 69.62±0.2 78.11±4.3 56.05±1.8 75.27±4.62 76.60±5.40 ResizeNet+5 AT - - 84.29±1.4 67.03±5.3 89.14±1.09 88.94±0.77 NST - - 84.87±1.7 70.11±1.0 89.07±0.15 89.46±0.28 SP - - 84.06±1.1 67.05±5.1 89.19±0.60 89.26±0.40 RKD - - 80.40±1.9 66.76±0.9 88.48±0.55 89.21±0.08 PKT - - 84.14±1.1 67.11±5.2 88.84±0.40 88.87±0.46 CC - - 83.54±1.2 69.26±1.0 89.46±1.39 88.64±1.40 MSA (ours) - - 87.03±0.9 75.03±1.6 90.55±0.73 91.34±0.85
  • 1
    Cross-dataset, within-species (human-to-human) and within-modality transfer: Train on human sEEG data from NICU or CHSZ dataset, and test on another human sEEG dataset. There are no channel differences between NICU and CHSZ datasets, thus the proposed ResizeNet+MSA approach was not performed in this setting.
  • 2
    Cross-dataset, within-species (human-to-human) and cross-modality transfer: Train on human iEEG/sEEG dataset, and test on human dataset with another modality, i.e., iEEG to sEEG, or sEEG to iEEG. Input discrepancies exist in this setting.
  • 3
    Cross-dataset, multi-species [(canine+human)-to-human] and multi-modality transfer: Canine iEEG and human iEEG/sEEG datasets were combined as the training set, and CHSZ was used as the test dataset. This setting used two species and two modality data to train the models, so that input discrepancies exist.
  • 4
    Baseline: Without using the proposed ResizeNet, the number of channels for both species was unified by eliminating the mismatching ones.
  • 5
    ResizeNet+: Utilize the proposed ResizeNet projection to unify the number of channels for different modalities.
  • 6
    Source Only: Train the model on all labeled data from the other dataset without employing any alignment strategy.
Table D1 shows the experiment results:
  1. 1.
    The cross-dataset within-species transfer performance was better than basic cross-species transfer without additional techniques.
  2. 2.
    ResizeNet+MSA demonstrated its effectiveness in cross-dataset cross-modality experiments.
  3. 3.
    Combining multi-species data in the training set and testing on a human dataset resulted in the best performance.
  4. 4.
    For the within-modality scenario, NICU and CHSZ datasets have the same number and placement of channels, so there were no data heterogeneous discrepancies. UDA approaches were not always effective. Compared to the cross-modality scenario, within-modality transfer achieved better performance, e.g., $85.17\%>84.37\%$ and $68.87\%>67.24\%$, indicating that cross-modality transfer (with data heterogeneous discrepancy) is indeed more challenging.
  5. 5.
    For the cross-modality scenario, iEEG and sEEG signals, collected from different devices, introduce significant data heterogeneity. While UDA approaches may be effective in some cases, the proposed ResizeNet framework consistently outperformed others. Notably, ResizeNet+MSA achieved the highest performance among all approaches.
  6. 6.
    For the multi-modality scenario, we further combined two-species datasets to build the training set. With the increasing number of training data, the performance improved, e.g., $84.37\%<86.33\%<87.05\%$. With the help of ResizeNet+MSA, the best performance was achieved on the CHSZ dataset, e.g., 91.34%, surpassing models trained on single-species single-modality data. This improvement likely stemmed from the increased training data volume and the proposed MSA, which effectively handles significant data heterogeneities.
Appendix E APPENDIX E: SEIZURE PREDICTION PERFORMANCE
Accurate forecasting of epileptic seizures greatly improves clinical epilepsy diagnosis and patient care. However, significant interspecies differences pose challenges for seizure prediction tasks. To evaluate the effectiveness of the proposed ResizeNet+MSA approach, extensive experiments were conducted for cross-species seizure prediction.
The Kaggle seizure prediction dataset from the seizure forecasting competition [60] was used, featuring epilepsy recordings from five canines and two humans undergoing extended iEEG monitoring. The data consist of 10-minute interictal and preictal clips. Substantial differences between canine and human recordings include sampling rates (400 Hz for canines, and 5000 Hz for humans) and electrode configurations (15 or 16 subdural electrodes for canines, and 15 or 24 intracranial channels for humans), reflecting cross-species characteristics similar to those in the seizure detection task. Additionally, the dataset exhibits a significant class imbalance between interictal and preictal trials. To address this imbalance, a subset of interictal samples was selected. For consistency with the seizure detection task, human iEEG signals were downsampled to 400 Hz, and each 10-minute clip was segmented into 10-second segments. Detailed characteristics of the dataset are summarized in Table E1.
Semi-supervised Canine-to-Human transfer experiments, with different portions of labeled trials, were conducted to evaluate three approaches: Within, Comb., and the proposed ResizeNet+MSA. Both Comb. and ResizeNet+MSA employed EA as a preprocessing step. The results, presented in Table E2, demonstrate that:
  1. 1.
    Comb. achieved superior performance compared to Within with an improvement of 6%, highlighting the benefits of incorporating data from another species.
  2. 2.
    The proposed ResizeNet+MSA achieved an additional 4.6% improvement over Comb., demonstrating its effectiveness in mitigating cross-species discrepancies and enabling better alignment across species.
Table E1: Characteristics of the five canine subjects and two human subjects in the Kaggle Prediction dataset.
Species ID # Channels Sampling rate (Hz) # Seizures # Preictal trials # Interictal trials # Total trials
Canine 1 16 400 22 1440 10,000 11,440
2 16 400 47 2520 10,000 12,520
3 16 400 104 4320 10,000 14,320
4 16 400 29 5820 10,000 15,820
5 15 400 19 1800 10,000 1,1800
Human 1 15 5,000 5 1,080 3,000 4,080
2 24 5,000 41 1,080 2,520 3,600
Table E2: Average semi-supervised cross-species seizure prediction transfer AUC (%) in Canine-to-Human seizure prediction. The best average performance of each task is marked in bold.
Approach Portion of labeled trials
5% 10% 15% 20% Avg.
Within 62.82±0.9 72.81±0.7 78.24±2.2 82.61±1.6 74.12
Comb. 76.62±1.0 82.11±0.4 83.80±1.4 84.59±1.8 81.78
ResizeNet+MSA 81.02±1.9 87.46±1.0 87.68±1.2 89.58±1.6 86.44